OpenStudy (riley):

Could someone please check my work for this problem. The automobile assembly plant you manage has a Cobb-Douglas production function given by: P = 10x^0.5*y^0.5 where P is the number of automobiles it produces per year, x is the number of employees, and y is the daily operating budget (in dollars). Assume that you maintain a constant work force of 130 workers and wish to increase production in order to meet a demand that is increasing by 50 automobiles per year. The current demand is 1000 automobiles per year. How fast should your daily operating budget be increasing?

OpenStudy (riley):

Typing out my work right now, give me just a second!!

OpenStudy (riley):

Typing out my work right now, give me just a second!!

OpenStudy (riley):

1000 = 5x^-0.5y^0.5 * dx/dt + 0.5y^-.510x^.5 * dy/dt = 5x^-.5*y^.5 * dp/dt + 5y^-.5*x^.5 * dy/dt =5(130)^-.5(77.44)^.5(50) + 5(77.44)^-.5(130^.5 * dy/dt =5(.09)(440) + 5(.11)(11.4) * dy/dt =192.95 + 6.48 * dy/dt =199.43*dy/dt 1000/199.43 = dy/dt dy/dt = 5.01

OpenStudy (riley):

I got y from this: P = 10x^0.5y^0.5 1000 = 10(130)^.5*y^.5 1000 = 10(11.4)y^.5 1000 = 114*y^.5 1000/114 =y^.5 8.8 = y^.5 (8.8)^2 = (y^.5)^2 77.44 = y